[Hallc_running] [Revised Logentry] RC Update

bduran at jlab.org bduran at jlab.org
Fri Sep 30 08:45:02 EDT 2022

Logentry Text:
Yesterday (September 29th) was the accelerator maintenance day. In hall C, we switched to the XEM2 experiments during the day shift. 

SHMS/HMS trigger setups were checked/modified according to the XEM2 experiment trigger requirements https://logbooks.jlab.org/entry/4048532, https://logbooks.jlab.org/entry/4048593. Data file locations were modified for the XEM2 experiments as well https://logbooks.jlab.org/entry/4048556. 

We started to set the magnets up for the first setting of the coincidence elastic running around 5pm. While ramping up to the MOL, HMS Q1 and Q2 threw an interlock. After visual inspection of the hall through the cameras, we successfully reset the interlock. Second attempt of the ramping up of these two magnets failed again but the resets worked. We managed to set the magnets up by increasing the current in increments rather than ramping it up to the MOL in one step. https://logbooks.jlab.org/entry/4048658 https://logbooks.jlab.org/entry/4048661

Around 6:45 pm we also had an issue with the HMS rotation while rotating it to the 19.52 degrees. The rotation stopped at 20.69 degrees when the drive controller faulted. Reached out to the engineer on call who needed an access to the hall to be able reset it. After the first reset the spectrometer did not want to cooperate with us. However, after the second reset we successfully rotated the HMS to the desired angle, 19.52 degrees. 

Beam was restored to the Hall C around 22:45. We did a quick harp scan and beam centering check with the Al hole target https://logbooks.jlab.org/entry/4048797 https://logbooks.jlab.org/entry/4048844 Unfortunately, the beam delivery was interrupted shortly after the restoration due the 1L12 trip and all RF in the North Linac. This was followed by the RSEP8A magnet over temperature fault. As a result of these issues, we currently do not have beam in hall C and we didn't receive any beam during the entire owl shift.

>From the 8 am MCC meeting --> RF in the NL recovered. Currently DC power is investigating the magnet. Beam will be restored upon understanding the issue with magnet and fixing it. 

We will have an RC meeting at 11 am today. Here is the link:



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4048909
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220930/50c93e70/attachment.html>

More information about the Hallc_running mailing list